When finding the volume of a box; (20-2x)(16-2x)(x) 1st term(20times16) = 320
3ed term(-2xtimes-2x)=+4x^2

How come the 2ed term is -72x??? I don't see the math. thank you

Answer by palanisamy(496)

(20-2x)(16-2x)(x)= (320-40x-32x+4x^2)*x
= (320-72x+4x^2)*x
= 320x-72x^2+4x^3